OpenStudy (anonymous):

F(x,y) = <-e^(-x)*ln(y), e^(-x)y^-1> compute a potential function for the conservative vector field F. I got e^(-x)ln(y) + C please help

OpenStudy (turingtest):

remember that C is a function of y possibly

OpenStudy (anonymous):

i integrated fx= -e^-x * ln y to get e^(-x) ln (y) + g(y) then I differentiated respect to y to get e^(-x) * 1/y + g'(y)

OpenStudy (turingtest):

\[\langle{\partial f\over\partial x},{\partial f\over\partial y}\rangle\]so\[f(x,y)=\int-e^{-x}\ln ydx=e^{-x}\ln y+g(y)\]since this is conservative we should be able to take the partial of this wrt y and get the other component

OpenStudy (anonymous):

i set fy to fy to get g'(y)= o

OpenStudy (turingtest):

\[{\partial f\over\partial y}=\frac1ye^{-x}+g'(y)\]so yeah, g'(y)=0 it would seem

OpenStudy (anonymous):

i integrated both sides to get g(y) = c...this is where I am confused

OpenStudy (turingtest):

so there will be a constant in the final answer, no big deal, that's normal

OpenStudy (anonymous):

part b of the problem state to evaluate from (0,e) to (0,e^2)

OpenStudy (turingtest):

so the function you got at the end is\[f(x,y)=e^{-x}\ln y+C\]and I suppose they want \(f(0,e)\) and \(f(0,e^2)\) so plug it in...

OpenStudy (anonymous):

how can i use the potential function to evaluate with C? e^-0 ln (e^2) - e^-0 ln (e) ??

OpenStudy (turingtest):

C will cancel if you subtract the two from each other

OpenStudy (turingtest):

so it doesn't matter

OpenStudy (anonymous):

oh ic...so it would be 2-1?

OpenStudy (anonymous):

=1

OpenStudy (turingtest):

I guess; never heard the term "evaluate" used for a function like that... but it seems like they want you to subtract them, which does give 1, yes

OpenStudy (anonymous):

thanks :)

OpenStudy (turingtest):

welcome!